Chemistry 96 – Math Component

Differential Calculus

Session C, Summer ’98 Lecture: M W 11:00AM to 1:00PM BH2760 Excel Workshop: F 11:00AM to 1:00PM Instructor: Ami Fischman Tutor: Pallavi Naidu

Topics: In this course we will introduce the notions of limits of functions, continuity, and derivatives (as tangents, velocities, etc.).

Textbook: We will use the book Calculus by James Stewart (ISBN 0-534-50072-2).

Homework: Homework will be assigned during each lecture, and collected during the following lecture. Math is not a spectator sport, and one does not learn it by listening alone. Make sure you start early on your homework, as some of it will be difficult, and starting early allows you to get help either from the tutor or the instructor. Homework is also a significant part of your grade, as itemized below.

Exams: There will be two exams: a midterm on 8/26 and a final on 9/21. The midterm will cover the material through section 2.5, and the final will cover all of chapter 2 (except for section 2.9, which we will skip). While understanding of the first 5 sections will be required, the final will emphasize the second half of the course more heavily.

Attendance & Quizzes: There will be a short quiz given at the beginning of each lecture (˜ 10 minutes). Additionally, your attendance during the Excel workshops (F 2-4) is mandatory. Pallavi will be holding tutorial sessions in the evenings (7-10, in griffin commons), and you are expected to show up sunday and tuesday nights (at least to sign in and show her your completed homework, if you have no questions about it). If you have time conflicts preventing you from making one of these meeting times, see Ami or Pallavi about possible special arrangements.

Grading: Your final grade for the math part of Chem 96 will be comprised of 4 parts: Homework: 15%, Midterm: 30%, Final: 40%, Attendance/Quizzes: 15%.

Contact Info: The best way to contact Ami is via email at: fischman@math.ucla.edu. Alternatively, you can leave him a note in his mailbox in MS6364 or try and catch him in his office, MS2344. The best way to contact Pallavi is via phone at (323)933-1757, or via email at pallavi@ucla.edu.

Announcements will be periodically posted to my web page at: http://www.math.ucla.edu/~fischman/ta/. Most handouts/solutions/etc from lecture will also be available here. If you are not comfortable with using the WWW or email, please ask either Ami or Pallavi and we will be happy to explain everything you need to know.

Tentative Syllabus: This is a draft of what I think we’ll be covering and when. The amount of material covered is likely to change – this is just a rough estimate.

Week Day  Section & Subject
1  M 8/10 2.1 The Tangent and Velocity Problems
W 8/12 2.2 The Limit of a Function
2  M 8/17 2.3 Calculating Limits
W 8/19 2.4 Continuity
3  M 8/24 2.5 Limits Involving Infinity
W 8/26 Midterm
4  M 8/31 2.6 Tangents, Velocities, and Other Rates of Change
W 9/2 2.7 Derivatives
5  M 9/7 Labor Day, no classes
W 9/9 2.8 The Derivative as a Function
6  M 9/14 2.10 The Relationship between f′ and f
W 9/16 Focus on Problem Solving (pp.186-189)
7  M 9/21 Final Exam